1.    A 10.0-liter flask contains 0.400 mole of H2, and 0.300 mole of N2 at 20°C.

(a)   What is the pressure, in atmospheres, inside the flask?

(b)   Calculate the mole fractions of the three gases?


1
Expert's answer
2021-10-14T04:05:40-0400

Moles of Nitrogen = 0.3

Moles of hydrogen = 0.4

Volume = 10 L

Temperature = 20 + 273 = 293 K

PV = nRT

Partial pressure of Nitrogen = 0.3×0.0821×208/10= 0.512atm

Partial pressure of hydrogen= 0.4×0.0821×208/10= 0.68atm

Total pressure = 0.512+0.68= 1.192atm